# Cryogenic source of atomic tritium for precision spectroscopy and neutrino-mass measurements

### v1

@@ Abstract
We propose a concept for a cryogenic source of atomic tritium at sub-kelvin temperatures and energies suitable for magnetic trapping. The source is based on the dissociation of solid molecular T2 films below 1 K by electrons from a pulsed RF discharge, a technique recently demonstrated for atomic hydrogen, combined with buffer-gas cooling and magnetic confinement. We analyze the key processes limiting the source performance, adsorption, spin exchange and recombination, and show that atomic tritium fluxes exceeding s-1 at kinetic energies of 100 mK can be achieved at the entrance of a magnetic trap. Such a source would enable Doppler-free two-photon 1S–2S spectroscopy in atomic tritium for high-precision measurements of the triton charge radius, providing a crucial benchmark for bound-state QED and improving the comparison between electronic, muonic, and scattering determinations of nuclear sizes in light systems. Beyond spectroscopy, an atomic tritium source avoids molecular final-state broadening in -decay and is therefore necessary for next-generation neutrino-mass measurements; combined with detector technologies such as sub-eV resolution quantum sensors or cyclotron radiation emission spectroscopy, it enables order-of-magnitude improvement compared to the current KATRIN sensitivity, reaching sensitivities below the inverted ordering regime. Additionally, the source can be used to generate a beam of low-field-seeking deuterium atoms for loading magnetic traps, an important benchmark before trapping tritium atoms and useful for precision spectroscopy.
@@ I Introduction
Studies of atomic hydrogen have played a central role in the development of modern physics. Because of its simplicity, many of its properties can be calculated from the first principles with extremely high accuracy, providing stringent tests of bound-state quantum electrodynamics (QED)Jentschura and Adkins (2022). The experimental possibility of reaching Bose–Einstein condensation stimulated extensive research on ultracold hydrogen atoms confined by superfluid helium films Silvera and Walraven (1986) or magnetic traps Greytak et al. (2000), where BEC was finally achieved in 1998 Fried et al. (1998). Renewed interest in such systems has recently emerged in connection with precision spectroscopy Brandt et al. (2022); Scheidegger and Merkt (2024); Amit et al. (2025), gravitational quantum states Vasiliev et al. (2019); Killian et al. (2024), and comparisons with antihydrogen to test the equivalence principle Ahmadi et al. (2018); Azevedo et al. (2023); Baker et al. (2025). Extending these studies to the heavier isotopes, deuterium and tritium, remains a major challenge, because of their stronger adsorption on helium surfaces and faster recombination. In the case of tritium, precision spectroscopy of the transition – would allow an accurate determination of the triton charge radius, providing an essential benchmark for few-body QED and consistency between the radii obtained from electronic, muonic, and scattering data. Such measurements would also contribute to the resolution of the persistent discrepancies among light-nuclear charge radii, which remain despite the recent convergence of some proton-radius determinations.
@@
Tritium is of particular interest because it is radioactive and decays by -emission into , an electron, and an antineutrino. Accurate measurements of the electron energy near the endpoint provide a direct determination of the effective electron-neutrino mass. This approach is followed by several major international collaborations, including KATRIN Aker et al. (2021), Project 8 Esfahani et al. (2017), QTNM Amad et al. (2025), and PTOLEMY Betti et al. (2019). The KATRIN experiment, which employs a windowless molecular source of T2 at 80 K, has recently established an upper limit of 0.45 eV for the neutrino mass, with a projected sensitivity of 0.3 eV in the final stage of its current measurement campaign Aker et al. (2025). These values remain above the expected lower bound of about 50 meV (9 meV) in the case of inverted mass hierarchy (normal hierarchy). One of the main limitations of the molecular approach is that part of the -decay energy is distributed into the ro-vibrational excitations of the final state of T2, which broadens the endpoint spectrum. This has motivated the development of sources of atomic tritium, where such molecular effects are absent, offering the possibility of reaching much higher precision in future direct neutrino-mass measurements.
@@
Techniques realized for magnetic trapping of H are now considered for holding large amounts of atomic D and T at temperatures 100 mK. Loading such a magnetic trap requires an efficient source of atoms with energies low enough for trapping with existing superconductive technologies. The dissociation of molecular H2, D2 and T2 can be carried out by thermal cracking Tschersich et al. (2008) or with a pulsed DC discharge combined with supersonic expansion Scheidegger et al. (2022). Although large fluxes and high degree of dissociation can be achieved with these methods, resulting atoms have very large energy and reducing it to the trappable values is a very difficult task.
@@
In this work, we propose the concept of a cryogenic source of atomic tritium. Dissociating T2 will be done below 1 K in a thin solid layer inside a dissociator chamber in a pulsed RF discharge similar to what is used to dissociate H2 Semakin et al. (2025). Electrons generated in tritium -decay at a mean energy of 5.7 keV propagate inside the solid layer and thus provide additional dissociation increasing the total atomic flux. Following H2 dissociation, the thermalization of the atoms usually takes place through collisions with the cold walls. However, this will not be feasible for tritium; therefore, instead buffer gas cooling with 4He or 3He vapor combined with radial magnetic compression is proposed. We evaluate that atomic fluxes of the order of 1015 atoms s-1 can be obtained at a gas temperature of 0.4 (0.2) K using 4He (3He) inside the gas transport tube. Finally, we present how evaporative cooling during gas transport towards the magnetic trap can be used for reaching even lower temperatures.
@@ II Properties of H, D and T
Since the pioneering experiment of Silvera and Walraven in 1980 Silvera and Walraven (1980a) in which spin-polarized H gas (H) was stabilized in a container lined with a superfluid helium film, a decade of intensive research was devoted to the experiments with high-field seeking (hfs) atoms (H) stabilized in a strong magnetic fields. Interaction with helium films was well understood and finally it became clear that the adsorption of the gas on the walls is the main obstacle and that the further progress towards BEC requires wall-free confinement in magnetic traps. Finally, BEC was reached at MIT in 1998 Fried et al. (1998) by evaporative cooling of low-field seeking (lfs) atoms (H) to K in a magnetic trap. Although the atoms were trapped and cooled down without contact with material walls, superfluid 4He coverage was needed to get a cold H beam with low enough temperature down to 100-200 mK for loading the trap.
Using similar techniques as those with H, several experiments succeeded in stabilizing atomic D inside a container lined with a superfluid helium film in zero Hayden and Hardy (1995); Arai et al. (1998) and strong magnetic fields Silvera and Walraven (1980b); Shinkoda et al. (1986). The densities and holding times achieved were substantially lower than those for H. The loading of atomic D into a magnetic trap was not successful Steinberger (2004). The only experimental work with atomic T was performed by Tjukanov in a zero magnetic field Tjukanoff et al. (1988). Small densities of gaseous T were detected by hyperfine resonance at temperatures above 8 K. However, no measurable signal below that was observed, even at around 1 K when the surfaces were coated with superfluid helium. Obviously, working with D and T turned out to be much more difficult. The reasons for that are related to the stronger interaction with the helium covered wall and faster recombination and relaxation, which we review below.
II.1 Interaction with the walls covered by superfluid helium film
Finding surfaces that would be most suitable for the confinement of atomic hydrogen was a key issue at the early stage of experimental work with cold H gas. Liquid helium, due to its smallest density and the formation of a superfluid film below 1 K, turned out to be the only choice to stabilize H and DSilvera and Walraven (1980a, b). The adsorption energy () of H on pure 4He and 3He-4He mixture films has been measured in numerous experiments through studies of the two-body recombination in the adsorbed phase (see Safonov et al. (2001) and references therein) and by direct measurement of the surface and bulk densities using high field ESRShinkoda and Hardy (1991). We present a summary of the experimental data and theoretical prediction of the main properties characterizing the interaction of hydrogen and its isotopes with the helium films of 4He and 3He in Table 1.
@@ II.1 Interaction with the walls covered by superfluid helium film
The adsorption energy for H on the 4He film is measured with the best accuracy K Safonov et al. (2001) and it can be seen that the surface coverage increases exponentially below 1 K. This sets a limit for the experimentally accessible temperature range down to 100 mK below which surface recombination leads to very large loss of atoms. In the mixture 3He-4He films at temperatures below 200-300 mK, 3He also fills the surface-bound (Andreev) states Pavloff and Treiner (9951). This leads to a substantial decrease in the adsorption energy of H on such mixture films down to 0.36 K Safonov et al. (2001). Due to a larger mass, D has a stronger interaction with the helium film, and the adsorption energy increases over 3 K on 4He Mosk et al. (2001). This limits the lowest experimental temperatures to around 300 mK provided that the recombination rate constant remains the same, which is actually not the case, as we shall see in the following. There are no experimental values for the adsorption energy of T. Following the trend between H and D and theoretical estimates, one may expect that it lies between 4 and 5 K for 4He and 1.5-2.5 K for films 3He.
The possibility of the atoms to penetrate into the bulk of the liquid helium film (sink) and adsorb on the solid substrate under the film was first considered theoretically. The chemical potential for H inside the liquid 4He (solvation energy) Es was evaluated as 36 KKurten and Ristig (1985); Saarela and Krotscheck (1993) and the solvation probability is negligible below 1 K. was predicted for D between 14 and 15 KKurten and Ristig (1985); Saarela and Krotscheck (1993). Because of the stronger binding, in order to keep the same recombination rate, the experimental temperature range can be increased by a factor of 3 compared to H. This goes above 1 K and the losses due to solvation can become substantial. Indeed, an exponential loss of D was observed in experiments of the UBC group Reynolds et al. (1991), and they reported a measurement of 13.6(6) K. For T, the predicted solvation energy K imposes even more serious problem, limiting upper operating temperature somewhere around 0.7 K. This is lower than the limit set by the adsorption energy, leading to the expectation that T cannot be stabilized in a container lined with the 4He helium film. Some chance remains if 3He-4He mixture films reduce for tritium as observed for hydrogen. At small concentrations of 3He, where the solvation energy may remain roughly the same, the adsorption energy may decrease by a factor of 3, leaving a possibility of experiments in a range between 0.3 and 0.7 K where the adsorbed phase density is not very large while the solvation probability is still low. To our knowledge, such experiments were not attempted.
@@
| H theor | 0.85Stwalley (1982) | 0.36Stwalley (1982) | 36Kurten and Ristig (1985); Saarela and Krotscheck (1993) | |
| H exp | 1.14(1) Safonov et al. (2001) | 0.39(1) Safonov et al. (2001) | 0.4(1)Safonov et al. (2001) | |
| D theor | 2.2(-0.7)Stwalley (1982) | 1.2Stwalley (1982) | 14Kurten and Ristig (1985), 15Saarela and Krotscheck (1993) | |
| D exp | 3.1(2)Mosk et al. (2001) | 13.6(6)Reynolds et al. (1991) | 30Hayden and Hardy (1995), 300Mayer and Seidel (1985) | |
| T theor | 3.2Stwalley (1982) | 1.9Stwalley (1982) | 6Kurten and Ristig (1985), 7.2Saarela and Krotscheck (1993) | |
| T exp | 10000Tjukanoff et al. (1988) |
@@ II.2 Surface recombination
Recombination of two hydrogen atoms to a molecule releases 55000 K of energy, being the main obstacle in experiments with hydrogen atoms at low temperatures. This process requires a third body, which can be another atom or a surface. Three-body processes that occur in the bulk gas or on the surface at very high densities are not considered in this work. Therefore, the two-body recombination on the surface of helium films is the main loss mechanism if the atoms are allowed to collide with the surface, as in experiments with the high-field seeking atomic states. The surface recombination rate is proportional to the collision rate of the atoms and to the effective cross-section of recombination which describes the probability of recombination in the collision. The corresponding rate constant can be expressed as where is the average thermal velocity in 2D, and is the recombination cross-length which may depend on temperature and magnetic field as well as on the hyperfine states of colliding atomsSilvera and Walraven (1986). The atoms in the so-called ”pure” high-field seeking hyperfine states (, in high field basis) for H and T and (, ) for D cannot recombine in two-body collisions. Therefore, the gas consisting of such states is extremely stable with respect to recombination until three-body processes start to play a role. At low densities, such doubly polarized gas decays via the first- and second-order processes of nuclear relaxation on the surface. For H, all rate surface recombination processes are well established and rate constants are accurately measured Silvera and Walraven (1986).
@@
The situation with the surface recombination for D and T is much less understood. Few experiments that attempted to work with D reported much faster recombination rates than those for H with maximum-obtained bulk densities at least an order of magnitude lower. These high rates could not be explained by the factor of 2.5-3 increase in adsorption energy (see Table 1). Discussions in refs. Arai et al. (1998); Mosk et al. (2001) considered a complicated behavior of the effective recombination cross-length dependent on the magnetic field and temperature, as well as the possibility of resonant recombination at certain values of B (Feshbach resonances). For the case of deuterium, the recombination cross-length may exceed the value for H by 2-3 orders of magnitude for a certain and . For T, the only experimental attempt Tjukanoff et al. (1988) did not succeed in stabilizing the T gas below 1 K. Solvation in the helium film and large recombination cross-section were suggested for the possible explanation. In conclusion, the superfluid helium film coverage, which successfully prevents surface recombination in the H gas, works poorly for D and most likely will not work for T.
II.3 Magnetic trapping of H, D, and T
Since surface recombination limits the cooling H states below 100 mK, a wall-free confinement method was suggested by HessHess (1986). It is based on trapping low field seekers () in a potential well created by a magnetic field. Cooling of the trapped gas is performed by evaporation of the atoms, which have enough energy to overcome the trapping barrier and carry away the energy. The remaining gas is thermalized in elastic two-body collisions. Successive trapping and cooling experiments were performed at MIT Hess et al. (1987) and the University of Amsterdam van Roijen et al. (1988) reaching BEC in 1998 Fried et al. (1998). Recently, a large magnetic trap for H was built and tested in the University of Turku Ahokas et al. (2022). The typical range of H densities confined in magnetic traps is between and cm-3 and minimum temperatures are several tens of K.
The trappable low field seeking states c (, in zero field basis) and d (, ) are higher in energy than two other hyperfine states a (, ) and b (, ). Relaxation to the lower energy states is threshold-less and does not vanish even at zero temperature. Spin exchange during two-body collisions is the fastest loss channel. It occurs in collisions where the ”mixed” state c takes part leading to a rapid depletion of the c state in the trap. The remaining ”pure” and doubly polarized state d is much more stable since spin-exchange does not work in this case. However, the dipolar interaction during collision of two d atoms may lead to spin flip and relaxation to untrapped states. This channel of dipolar relaxation is the main loss mechanism that limits the lifetime and maximum densities of the trapped gas. This is a second order, in density, process with the characteristic decay time inversely proportional to the gas density. For H, the two-body dipolar relaxation rate constant is cm3/s Stoof et al. (1988), which implies a lifetime of s at the density cm-3.
The stability of trapped T gas is basically determined by similar relaxation processes as in H. Both atoms are bosons and have a similar hyperfine structure. However, the three times larger mass has a strong effect. As recently calculatedElliott and Jones (2025), the low-energy cross sections for T-T collisions are substantially larger than those for H-H. This concerns both elastic and inelastic collisions. The reason is that there appears a scattering resonance in the collisions of the atoms with the hypothetical mass slightly above 3Elliott and Jones (2025). Two T atoms are very close to the formation of the bound dimer state. The s-wave scattering length is very large and negative Å. This increases the elastic collision rate and enhances the efficiency of evaporative cooling. However, inelastic dipole relaxation is increased as well. For the relaxation d-d of two T atoms in a magnetic field of 1 mT and temperature of 1 mK, typical for H trapping, the increase reaches a factor of compared to that for H-H collisions. The dipolar rate for T-T collisions has a much stronger temperature dependence, and the difference with H decreases to a factor of for T=100 mK. This leads to an increase in the loss rate and a decrease in the life-time of the trapped T gas by the same factors at equal density. To our knowledge, magnetic trapping of T has not been attempted to date, and the theoretical predictions above are still waiting for experimental confirmation.
The situation is very different for fermionic D. Due to the Pauli principle, identical D atoms avoid approaching each other, and s-wave scattering is forbidden. This would allow substantially larger densities and very high stability of the trapped gasKoelman et al. (1988). However, the only attempt to trap the D gas performed by the MIT group using the same dissociation and trapping techniques as for HSteinberger (2004) was unsuccessful. The walls of their trapping cell were covered with superfluid helium film. Although some flux of D entering the trap was detected, no signal was seen after the dissociator was switched off. The large adsorption energy on the helium film and fast recombination rate on the surface of the trapping cell were suggested as a possible explanation for their failure to load the trap.
II.4 Scattering of H isotopes on each other and on He

### v2

# Cryogenic source of atomic tritium for neutrino-mass measurements and precision spectroscopy

@@ Abstract
We propose a concept for a cryogenic source of atomic tritium at sub-Kelvin temperatures and energies suitable for magnetic trapping. The source is based on the dissociation of solid molecular T2 films below 1 K by electrons from a pulsed RF discharge, a technique recently demonstrated for atomic hydrogen, combined with buffer-gas cooling and magnetic confinement. We analyze the key processes limiting the source performance, adsorption, spin exchange and recombination, and show that atomic tritium fluxes exceeding s-1 at kinetic energies of 100 mK can be achieved at the entrance to the magnetic trap. Such a source would enable Doppler-free two-photon 1S–2S spectroscopy in atomic tritium for high-precision measurements of the triton charge radius, providing a crucial benchmark for bound-state QED and improving the comparison between electronic, muonic, and scattering determinations of nuclear sizes in light systems. Beyond spectroscopy, an atomic tritium source avoids molecular final-state broadening in the -decay and is therefore necessary for next-generation neutrino-mass measurements; combined with detector technologies such as sub-eV resolution quantum sensors or cyclotron radiation emission spectroscopy, it enables an order-of-magnitude improvement compared to the current best experimental limit. Additionally, the source can be used to generate a beam of low-field-seeking deuterium atoms for loading magnetic traps, an important benchmark before trapping tritium atoms, which is useful for precision spectroscopy.
@@ I Introduction
Studies of atomic hydrogen have played a central role in the development of modern physics. Due to its simplicity, many of its properties can be calculated from the first principles with extremely high accuracy, providing stringent tests of bound-state quantum electrodynamics (QED) [1]. The experimental possibility of reaching Bose–Einstein condensation stimulated extensive research on ultracold hydrogen atoms confined by superfluid helium films [2] or magnetic traps [3], where BEC was finally achieved in 1998 [4]. Renewed interest in such systems has recently emerged in connection with precision spectroscopy [5, 6, 7], gravitational quantum states [8, 9], and comparisons with antihydrogen to test the equivalence principle [10, 11, 12]. Extending these studies to the heavier isotopes, deuterium and tritium, remains a major challenge, because of their stronger adsorption on helium surfaces and faster recombination. In the case of tritium, precision spectroscopy of the transition – would allow an accurate determination of the triton charge radius, providing an essential benchmark for few-body QED and consistency between the radii obtained from electronic, muonic, and scattering data [13]. Such measurements would also contribute to the resolution of persistent discrepancies among light-nuclear charge radii, which remain despite of recent convergence of some proton-radius determinations [14].
@@
Tritium is of particular interest because it is radioactive and decays by -emission into , an electron, and an antineutrino. Accurate measurements of the electron energy near the endpoint provide a direct determination of the effective electron-neutrino mass. This approach is pursued by several international collaborations, including the Karlsruhe Tritium Neutrino Experiment (KATRIN) [15], Project 8 [16], QTNM [17] and PTOLEMY [18]. The KATRIN experiment, which employs a windowless molecular source of T2 at 80 K, has recently established an upper limit of 0.45 eV for the neutrino mass, with a projected sensitivity of 0.3 eV (both at 90 confidence level) in the final stage of its current measurement campaign [19]. These values remain above the expected lower bound of about 50 meV (9 meV) in the case of an inverted mass hierarchy (normal hierarchy). The normal and inverted hierarchies denote the two possible orderings of the neutrino mass eigenstates , , and consistent with oscillation data. In the normal (inverted) ordering one has (), implying different minimal values for the effective electron-neutrino mass accessible to direct kinematic searches [20]. One of the main limitations of the molecular approach is that part of the -decay energy is distributed into the ro-vibrational excitations of the final state of T2, which broadens the endpoint spectrum [21]. This has motivated the development of sources of atomic tritium, where such molecular effects are absent, offering the possibility of reaching much higher precision in future direct neutrino-mass measurements.
@@
Techniques realized for magnetic trapping of H are now considered for holding large amounts of atomic D and T at temperatures 100 mK. Loading such a magnetic trap requires an efficient source of atoms with energies low enough for trapping with existing superconductive magnet technologies. The dissociation of molecular H2, D2 and T2 can be carried out in RF discharge [22], by thermal cracking [23] or with a pulsed DC discharge combined with supersonic expansion [24]. Although large fluxes and high degree of dissociation can be achieved with these methods, resulting atoms have very large energy and reducing it to the trappable values is a very difficult task.
@@
In this work, we propose a concept of a cryogenic source of atomic tritium. Dissociating T2 will be done below 1 K in a thin solid layer inside a dissociator chamber in a pulsed RF discharge similar to what is used to dissociate H2 [25]. Electrons generated in tritium -decay at a mean energy of 5.7 keV propagate inside the solid layer and thus provide additional dissociation increasing the total atomic flux. For the H source, following dissociation, the thermalization of the atoms usually takes place through collisions with the cold walls. However, this will not be feasible for tritium. Therefore, buffer gas cooling with 4He or 3He vapor combined with radial magnetic confinement is proposed in this work. This method was utilized to cool and trap several atomic and molecular species after laser ablation from the solid target in J. Doyle’s group (see [26] for a review). We evaluate that atomic fluxes of the order of 1015 T atoms s-1 can be obtained at gas temperature of (0.2) K using 4He (3He) inside the gas transport tube.
@@ II Properties of H, D and T
Since the pioneering experiment of Silvera and Walraven in 1980 [2] in which spin-polarized H gas (H) was stabilized in a container lined with a superfluid helium film, a decade of intensive research was devoted to the experiments with high-field-seeking (hfs) atoms (H) stabilized in strong magnetic fields (see the hyperfine level diagram in Fig. 1). The interaction with helium films was well understood and finally it became clear that the adsorption of the gas on the walls is the main obstacle and that further progress towards BEC requires wall-free confinement in magnetic traps. Finally, BEC was reached at MIT in 1998 [4] by evaporative cooling of low-field seeking (lfs) atoms (H) to K in a magnetic trap. Although the atoms were trapped and cooled without contact with the material walls, a superfluid 4He coverage was needed to provide a cold H beam with a sufficiently low temperature down to 100-200 mK for loading the trap.
Using techniques similar to those with H, several experiments succeeded in stabilizing atomic D inside a container lined with a superfluid helium film in zero [27, 28] and strong magnetic fields [29, 30]. The densities and holding times achieved were substantially lower than those for H. The loading of atomic D into a magnetic trap was not successful [31]. The only experimental work with atomic T was performed by Tjukanov in zero magnetic field [32]. Small densities of gaseous T were detected by the hyperfine resonance at temperatures greater than 8 K. However, no measurable signal below that was observed, even at around 1 K when the surfaces were coated with superfluid helium. Working with D and T turned out to be much more difficult. The reasons for this are related to the stronger interaction with the helium covered wall and faster recombination and relaxation, which we review below.
II.1 Interaction with the walls covered by a superfluid helium film
Finding surfaces that would be most suitable for the confinement of atomic hydrogen was a key issue at the early stage of experimental work with cold H gas. Liquid helium, due to its smallest density and the formation of a superfluid film below 1 K, turned out to be the only choice to stabilize H and D [2, 29]. The adsorption energy () of H on films of a pure mixture of 4He and 3He-4He has been measured in numerous experiments through studies of a two-body recombination in the adsorbed phase (see [33] and references therein) and by direct measurement of the surface and bulk densities using high field ESR [34]. We present a summary of the experimental data and theoretical predictions of the main properties characterizing the interaction of hydrogen and its isotopes with the helium films of 4He and 3He in Table 1.
@@ II.1 Interaction with the walls covered by a superfluid helium film
The adsorption energy for H on the 4He film is measured with the best accuracy K [33] and it can be seen that the surface coverage increases exponentially below 1 K. This sets a limit for the experimentally accessible temperature range down to 100 mK below which surface recombination leads to a very large loss of atoms. In the 3He-4He mixture films at temperatures below 200-300 mK, 3He also fills the surface-bound (Andreev) states [35]. This leads to a substantial decrease in the adsorption energy of H on such mixture films down to 0.36 K [33]. Due to a larger mass, D has a stronger interaction with the helium film, and the adsorption energy increases over 3 K on 4He [36]. This limits the lowest experimental temperatures to around 300 mK provided that the recombination rate constant would remain the same, which is actually not the case, as we shall see in the following. To our knowledge, there are no experimental values for the adsorption energy of T. Following the trend between H and D and theoretical estimates, one may expect that it lies between 4 and 5 K for 4He and 1.5-2.5 K for films 3He.
Some chance remains that 3He-4He mixture films reduce for tritium, as observed for hydrogen. At small concentrations of 3He, where the solvation energy may remain roughly the same, the adsorption energy can decrease by a factor of 3, leaving the possibility of experiments in a range between 0.3 and 0.7 K where the adsorbed phase density is not very large while the solvation probability is still low. To the best of our knowledge, such experiments were not attempted.
@@
| H theor | 0.85 [37], 1.00 [38] | 0.36 [37] | 36 [39, 40] | |
| H exp | 1.14(1) [33] | 0.39(1) [33] | 0.4(1) [33] | |
| D theor | 2.2 [37], 2.54 [38] | 1.2 [37] | 14 [39], 15[40] | |
| D exp | 3.1(2) [36] | 13.6(6) [41] | 30 [27], 300 [42] | |
| T theor | 3.2 [37], 4.77 [38] | 1.9 [37] | 6 [39], 7.2 [40] | |
| T exp | 10000 [32] |
@@ II.2 Surface recombination
Recombination of two hydrogen atoms to a molecule releases 52000 K of energy, being the main obstacle in experiments with hydrogen atoms at low temperatures. This process requires a third body, which can be another atom or a surface. Three-body processes that occur in the bulk gas or on the surface at very high densities are not considered in this work. Therefore, the two-body recombination on the surface of helium films is the main loss mechanism if the atoms are allowed to collide with the surface, as in experiments with the high-field-seeking atomic states. The surface recombination rate is proportional to the collision rate of the atoms and to the effective cross-section of the recombination which describes the probability of recombination in the collision. The corresponding rate constant can be expressed as where is the average thermal velocity in 2D, and is the recombination cross-length which may depend on temperature and magnetic field as well as on the hyperfine states of colliding atoms [43]. The atoms in the so-called ”pure” high-field seeking hyperfine states (, ) for H and T and (, ) for D (see Fig. 1) cannot recombine in two-body collisions. Therefore, the gas consisting of such states is extremely stable with respect to recombination until three-body processes start to play a role. At low densities, such doubly polarized gas decays via the first- and second-order processes of nuclear relaxation on the surface. For H, all rate surface recombination processes are well established and rate constants are accurately measured [43].
@@
The situation with the surface recombination for D and T is much less understood. Few experiments that attempted to work with D reported much faster recombination rates than those for H with maximum-obtained bulk densities at least an order of magnitude lower. These high rates could not be explained by the factor of 2.5-3 increase in the adsorption energy (see Table 1). Discussions in refs. [28, 36] considered a complicated behavior of the effective recombination cross-length dependent on the magnetic field and temperature, as well as the possibility of resonant recombination at certain values of B (Feshbach resonances). For the case of deuterium, the recombination cross-length can exceed the value for H by 2-3 orders of magnitude for a certain and . For T, the only experimental attempt [32] did not succeed in stabilizing the T gas below 1 K. Solvation in the helium film and a large recombination cross-section were proposed for the possible explanation. In conclusion, the superfluid helium film coverage, which successfully prevents surface recombination in the H gas, works poorly for D and most likely will not work for T [41].
II.3 Scattering of H isotopes on each other and on He